OpenStudy (dollbeautyx):

1. Find the equation of the line written in point-slope form that passes through the point (1, –6) and is parallel to the line y = –8x + 3

OpenStudy (dollbeautyx):

@iGreen do you know how to do this.

OpenStudy (skullpatrol):

Do you know what "point-slope" form looks like?

OpenStudy (dollbeautyx):

y-y= m(x-x)?

OpenStudy (skullpatrol):

The formula $$y - y_1 = m(x - x_1)$$ is usually described as the 'point-slope form' for the equation of a line.

OpenStudy (dollbeautyx):

After knowing the formula, How do I write the equation?

OpenStudy (skullpatrol):

good question :-) you need to find a point x sub 1 and y sub 1 ON the line

OpenStudy (dollbeautyx):

So can it be any coordinate that passes through (1,-6)?

OpenStudy (skullpatrol):

yes, it can even be (1, -6)

OpenStudy (skullpatrol):